OpenStudy (anonymous):

The scale drawing for a house is 1.5 inches:12 feet. If one side of the garage is 40 feet, how long would that be in the drawing?

OpenStudy (anonymous):

Also, you draw a name at random from a hat holding the names of 6 girls and 8 boys. What are the odds in favor of choosing a girl? Please answer both. ;~; Thanks.

OpenStudy (anonymous):

okay so every 1.5 inches is 12 feet right? how many 12 feets are there in 40 feet? divide 40 by 12 = 3.33 times 1.5 by 3.33 = 4.99 inches scale drawing should be 4.99 inches

OpenStudy (anonymous):

6 girls and 8 boys is 14 kids probability is : \[\frac{ number you want }{ total number } = \frac{ 6 }{ 14 } = \frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

or 0.428 as a decimal

OpenStudy (anonymous):

Thank you so much! :')